Exercise 10-5 Working Backwards from Labor Variances [LO10-2]

The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

  Job Standard Hours   Standard Rate   Standard Cost  
  Motor tune-up 2.5         $9.00         $22.50       

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 50 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

  Labor rate variance $ 87 F
  Labor spending variance $ 93 U
Required:
1. Determine the number of actual labor-hours spent on tune-ups during the week.

     

2.

Determine the actual hourly rate of pay for tune-ups last week.
Answer #1

Labor efficiency variance = 93U - 87F = 180 U

Labor efficiency variance = (SH-AH) *SR

(2.5*50 - AH) * 9 =-180

125 - AH = - 20

ACTUAL HOURS = 145

Labpr rate variance = (SR-AR) *AH

(9 - AR) * 145 = 87

(9-AR) = 0.6

ACTUAL HOURLY RATE = 8.4